OpenStudy (anonymous):

Mark and Nancy are making punch for a party and are using two types of juice: pineapple and papaya. The total number of quarts must be more than 9 but less than 15. They should use at least 4 quarts of each type. We want to find out how many quarts of each juice they should buy. List any 3 ordered pair solutions for the system of equations which represent this problem.

OpenStudy (dumbcow):

set up inequalities Let x be pineapple juice, y be papaya juice x>=4 y >=4 9 < x+y < 15 --> 9-x < y < 15 -x just pick (x,y) combinations that satisfy these conditions

OpenStudy (anonymous):

Would these work ? 4 + 6 = 10 : (4 , 6) 4 + 7 = 11 : (4 , 7) 4 + 8 = 12 : (4 , 8) 4 + 9 = 13 : (4 , 9)

OpenStudy (dumbcow):

yep all of those would work